# Linear Discriminant Regularized Regression

### v1

@@ Abstract
Linear Discriminant Analysis (LDA) is an important classification approach. Its simple linear form makes it easy to interpret and it is capable to handle multi-class responses. It is closely related to other classical multivariate statistical techniques, such as Fisher’s discriminant analysis, canonical correlation analysis and linear regression. In this paper we strengthen its connection to multivariate response regression by characterizing the explicit relationship between the discriminant directions and the regression coefficient matrix. This key characterization leads to a new regression-based multi-class classification procedure that is flexible enough to deploy any existing structured, regularized, and even non-parametric, regression methods. Moreover, our new formulation is generically easy to analyze compared to existing regression-based LDA procedures. In particular, we provide complete theoretical guarantees for using the widely used $\ell_{1}$-regularization that has not yet been fully analyzed in the LDA context. Our theoretical findings are corroborated by extensive simulation studies and real data analysis.
@@ 1 Introduction
Here $\{{\bm{e}}_{1},\ldots,{\bm{e}}_{L}\}$ are the canonical basis vectors in $\mathbb{R}^{L}$ and the within-class covariance matrix $\Sigma_{w}$ is assumed to be strictly positive definite. For any $i\in[n]$, we say that observation $X_{i}$ belongs to class $\ell\in[L]$ if its label $Y_{i}={\bm{e}}_{\ell}$, which one may simply write as $Y_{i}=\ell$. Our insistence on the perhaps unusual notation ${\bm{e}}_{\ell}$ will become transparent later when we formulate the estimation of the optimal hyperplane as a multivariate response regression problem. Since we can always subtract the marginal mean of $X$, we assume
$\mathbb{E}(X)=0_{p}$ (1.2)
for simplicity, and furthermore that the class probabilities
$\pi_{\ell}=\mathbb{P}\{Y={\bm{e}}_{\ell}\}$ (1.3)
are strictly positive for all $\ell\in[L]$. The Bayes rule under model (1.1) and (1.3) is to classify a new feature $x\in\mathbb{R}^{p}$ as
@@
Few approaches consider (ii) but notable exceptions are Abramovich and Pensky (2019); Nibbering and Hastie (2022). We recall various approaches to deal with (i). Write
$\pi=(\pi_{1},\ldots,\pi_{L})^{\top},\qquad D_{\pi}={\rm diag}(\pi),\qquad{\mathbb{M}}=(\mu_{1},\ldots,\mu_{L})\in\mathbb{R}^{p\times L}$ (1.5)
@@
$\Sigma={\rm Cov}(X)=\Sigma_{w}+{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}.$
We call $\Sigma_{w}$ the within-class covariance matrix and $\Sigma_{b}:=\Sigma-\Sigma_{w}={\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}$ the between-class covariance matrix. The latter due to (1.2) has rank less than $L$.
In high-dimensional settings $p>n$, Cai and Liu (2011) assume that the direction $\Sigma_{w}^{-1}(\mu_{2}-\mu_{1})$ is sparse in the binary case $L=2$, and propose a Dantzig-type procedure to estimate this direction (see, also, Cai and Zhang (2019) for extension to $L>2$). It isn’t always reasonable to assume that such directions are sparse and the procedure is less attractive compared to regression-based approaches discussed later. Witten and Tibshirani (2011) use the fact that the Bayes rule (1.4) is equivalent with Fisher’s discriminant rule
@@
$\begin{split}F_{k}:=\operatorname*{arg\,max}_{\beta\in\mathbb{R}^{p}}&~{}\beta^{\top}\Sigma_{b}\beta\\ \text{such that}&\quad\beta^{\top}\Sigma_{w}\beta=1,\quad\beta^{\top}\Sigma_{w}F_{i}=0,~{}~{}\forall~{}i<k.\end{split}$ (1.7)
Witten and Tibshirani (2011) develop methodology tailored to the high-dimensional scenarios with a diagonal matrix $\Sigma_{w}$ and sparse columns of $F$. However, the implicit formulation (1.7) of the matrix $F$ makes it extremely difficult to obtain the global solution and to analyze the theoretical properties of the regularized estimator, hence the resulting classifier. More closely related to our approach, introduced in the next paragraph, is optimal scoring, which dates back to De Leeuw et al. (1976); Young et al. (1978) and regains attention in the LDA setting, see Hastie et al. (1995, 1994); Clemmensen et al. (2011); Gaynanova (2020) and the references therein.
By writing the response matrix and feature matrix as
${\mathbb{Y}}=(Y_{1},\ldots,Y_{n})^{\top}\in\{0,1\}^{n\times L},\quad{\mathbb{X}}=(X_{1},\ldots,X_{n})^{\top}\in\mathbb{R}^{n\times p},$ (1.8)
@@
$\begin{split}\min_{\ \theta\in\mathbb{R}^{L},\ \beta\in\mathbb{R}^{p}}&~{}\|{\mathbb{Y}}\theta-{\mathbb{X}}\beta\|_{2}^{2}\\ \text{such that }&~{}\theta^{\top}{\mathbb{Y}}^{\top}{\mathbb{Y}}\theta=n,\quad\theta^{\top}{\mathbb{Y}}^{\top}{\mathbb{Y}}\widehat{\theta}_{i}=0,~{}\forall~{}i<k.\end{split}$ (1.9)
@@
Here $\widehat{\theta}_{i}$ is the solution to (1.9) in the $i$th iteration. The minimization problem (1.9) can be done via solving a generalized eigenvalue problem, see Hastie et al. (1995). Moreover, Hastie et al. (1995) also show that the solutions $\widehat{\beta}_{1},\ldots,\widehat{\beta}_{K}$ from (1.9) can be computed from canonical correlation analysis (CCA), and are further related, via CCA, with the matrix $F$ obtained from the sample analogue of (1.7) in the low-dimensional case ($p<n$). As a result, classification can be based on the dimension reduction directions $\widehat{\beta}_{1},\ldots,\widehat{\beta}_{K}$.
Since optimal scoring is a regression-based approach, as pointed out in Hastie et al. (1995, 1994), it is often easier to compute and more amenable to extensions to regularized and non-parametric regression as well as model selection than canonical correlation analysis, Fisher’s discriminant rule and the procedure of Cai and Zhang (2019).
In the high-dimensional setting, Hastie et al. (1995) show that the global solution to a penalized version of (1.9), that is, for $1\leq k\leq K<L$,
@@
$\begin{split}\min_{\theta\in\mathbb{R}^{L},\ \beta\in\mathbb{R}^{p}}&~{}\|{\mathbb{Y}}\theta-{\mathbb{X}}\beta\|_{2}^{2}+\beta^{\top}\Omega\beta\\ \text{such that}&~{}\theta^{\top}{\mathbb{Y}}^{\top}{\mathbb{Y}}\theta=n,\quad\theta^{\top}{\mathbb{Y}}^{\top}{\mathbb{Y}}\widehat{\theta}_{i}=0,~{}\forall~{}i<k,\end{split}$ (1.10)
@@
can be computed for any $p\times p$ symmetric, positive definite matrix $\Omega$. A possible choice of $\Omega=\lambda\mathbb{I}_{p}$ for some $\lambda>0$ leads to the ridge penalty (Campbell, 1980; Friedman, 1989). In a more recent work (Gaynanova, 2020), the matrix formulation of (1.10)
$\begin{split}\min_{\Theta,~{}B}&~{}\|{\mathbb{Y}}\Theta-{\mathbb{X}}B\|_{F}^{2}+\text{pen}(B)\\ \text{such that }&~{}\Theta^{\top}{\mathbb{Y}}^{\top}{\mathbb{Y}}\Theta=n\mathbb{I}_{K},\quad\Theta^{\top}{\mathbb{Y}}^{\top}{\mathbb{Y}}1_{n}=0_{K}.\end{split}$ (1.11)
@@
is considered, and it is proved that the global solution $\widehat{B}\in\mathbb{R}^{p\times K}$ to (1.11) can be computed for any penalty $\text{pen}(B)$ that satisfies $\text{pen}(BR)=\text{pen}(B)$ for any orthogonal $K\times K$ matrix $R$. This allows for the penalty $\|B\|_{1,2}=\sum_{j=1}^{p}\|B_{j\cdot}\|_{2}$ used in the group-lasso, but excludes the familiar $\|B\|_{1}=\sum_{j,k}|B_{jk}|$ used in the lasso regression as well as its extensions such as the elastic net penalty. That being said, for penalties that are not orthogonal invariant, there is no guarantee of computing the global solution to (1.11), let alone any theoretical property of the resulting estimator.
@@
Key to our approach is the following reformulation of the problem. We show in Lemma 1 of Section 2 that the discriminant directional matrix under model (1.1) and (1.3)
$B^{*}=\Sigma_{w}^{-1}{\mathbb{M}}\in\mathbb{R}^{p\times L}$ (1.12)
@@
$B:=\Sigma^{-1}\Sigma_{XY}:=[{\rm Cov}(X)]^{-1}{\rm Cov}(X,Y)$ (1.13)
@@
$B^{*}$ $=BH^{-1}$
for some invertible $L\times L$ matrix $H$. This connection is novel and distinguishes our approach from the current literature. Although the fact that $B$ and $B^{*}$ share the same span is relatively well-known (see, for instance, Hastie et al. (1995); Ye (2007); Lee and Kim (2015); Nie et al. (2022) and the references therein), their closed-form connection has not been discovered and a thorough theoretical investigation on the corresponding classifier has not been conducted, to the best of our knowledge. Specifically, we propose to estimate the discriminant matrix $B^{*}$ by estimating the regression matrix $B$ first
$\widehat{B}=\operatorname*{arg\,min}_{B\in\mathbb{R}^{p\times L}}~{}{1\over n}\|{\mathbb{Y}}-{\mathbb{X}}B\|_{F}^{2}+\text{pen}(B),$ (1.14)
and then estimating $H^{-1}$ based on $\widehat{B}$. The advantage of our approach is three-fold:
@@
    We can easily tap into the large literature on estimation of the high-dimensional regression matrix $B\in\mathbb{R}^{p\times L}$.
@@
    After estimating $B$, we only need to estimate the inverse of an $L\times L$ matrix rather than that of a $p\times p$ matrix $\Sigma_{w}$ as in the classical LDA context.
@@
    We are able to quantify the estimation error of both $B$ and $B^{*}$ as well as to provide theoretical guarantees for the resulting classifier.
While various penalties are feasible in our approach, more specifically in (1.14), we focus on a lasso-type estimator of $B$ in this paper. This also contrasts our approach with Cai and Zhang (2019) and Witten and Tibshirani (2011) which assume sparsity of $B^{*}$ and $F$, respectively. Our earlier work Bing and Wegkamp (2023) considers a multi-class classification setting with high-dimensional features $X\in\mathbb{R}^{p}$ that follow a factor model $X=AZ+W$ with unobserved (latent) features $Z\in\mathbb{R}^{r}$ and unobserved noise $W$, that is independent of both $Z$ and the label $Y$ with $r<n<p$. The procedure is entirely different than the one proposed here. Future work will be devoted on how to adapt the procedure in this paper to this low-rank setting and its analysis.
Under sparsity of the regression matrix $B$, we carefully analyze the properties of $\widehat{B}$ in Section 3.1, the estimation of $B^{*}$ via estimating $H^{-1}$ in Section 3.2 as well as the rate of convergence for the excess misclassification risk of the corresponding classification rule in Section 3.3. Computationally, after computing $\widehat{B}$ from (1.14) we offer an alternative way in Section 2.3 to estimate the Bayes rule that performs Fisher’s discriminant analysis on low-dimensional transformations $\widehat{B}^{\top}x\in\mathbb{R}^{L}$ of the original features $x\in\mathbb{R}^{p}$. This leads to a practically useful method, which extracts fewer discriminant directions than those in $B^{*}$, for both downstream analysis and visualization. Finally, we provide a thorough simulation study in Section 4 and we report our findings in several real data studies of Section 5.

### v2

@@ Abstract
Linear Discriminant Analysis (LDA) is an important classification approach. Its simple linear form makes it easy to interpret and it is capable to handle multi-class responses. It is closely related to other classical multivariate statistical techniques, such as Fisher’s discriminant analysis, canonical correlation analysis and linear regression. In this paper we strengthen its connection to multivariate response regression by characterizing the explicit relationship between the discriminant directions and the regression coefficient matrix. This key characterization leads to a new regression-based multi-class classification procedure that is flexible enough to deploy any existing structured, regularized, and even non-parametric, regression methods. Moreover, our new formulation is amenable to analysis: we establish a general strategy of analyzing the excess misclassification risk of the proposed classifier for all aforementioned regression techniques. As applications, we provide complete theoretical guarantees for using the widely used $\ell_{1}$-regularization as well as for using the reduced-rank regression, neither of which has yet been fully analyzed in the LDA context. Our theoretical findings are corroborated by extensive simulation studies and real data analysis.
@@ 1 Introduction
Here $\{{\bm{e}}_{1},\ldots,{\bm{e}}_{L}\}$ are the canonical basis vectors in $\mathbb{R}^{L}$ and the within-class covariance matrix $\Sigma_{w}$ is assumed to be strictly positive definite. For any $i\in[n]$, we say that observation $X_{i}$ belongs to class $\ell\in[L]$ if its label $Y_{i}={\bm{e}}_{\ell}$, which one may simply write as $Y_{i}=\ell$. Our insistence on the perhaps unusual notation ${\bm{e}}_{\ell}$ will become transparent later when we formulate the estimation of the optimal hyperplane as a multivariate response regression problem. Since we can always subtract the marginal mean of $X$, we assume $\mathbb{E}(X)=0_{p}$ throughout this paper mainly for simplicity, and furthermore that the class probabilities $\pi_{\ell}=\mathbb{P}\{Y={\bm{e}}_{\ell}\}$ are strictly positive for all $\ell\in[L]$. The Bayes rule under model (1.1) is to classify a new feature $x\in\mathbb{R}^{p}$ as
@@
Few approaches consider (ii) but notable exceptions are Abramovich and Pensky (2019); Nibbering and Hastie (2022). We recall various approaches to deal with (i). We will use the following notation. Write
$\pi=(\pi_{1},\ldots,\pi_{L})^{\top},\quad D_{\pi}={\rm diag}(\pi)\in\mathbb{R}^{L\times L},\quad{M}=(\mu_{1},\ldots,\mu_{L})\in\mathbb{R}^{p\times L}$ (1.3)
@@
$\Sigma={\rm Cov}(X)=\Sigma_{w}+{M}D_{\pi}{M}^{\top}.$ (1.4)
We call $\Sigma_{w}$ the within-class covariance matrix and $\Sigma_{b}:=\Sigma-\Sigma_{w}={M}D_{\pi}{M}^{\top}$ the between-class covariance matrix. The latter has rank less than $L$ due to $\mathbb{E}(X)=0_{p}$.
In high-dimensional settings $p>n$, Cai and Liu (2011) assume that the direction $\Sigma_{w}^{-1}(\mu_{2}-\mu_{1})$ is sparse in the binary case $L=2$, and propose a Dantzig-type procedure by approximately solving the equation $\Sigma_{w}\beta^{*}=\mu_{2}-\mu_{1}$ for $\beta^{*}$ to estimate this direction. Other procedures based on the same or similar equations include, for instance, Qiao et al. (2009); Shao et al. (2011); Fan et al. (2012) and Tibshirani et al. (2002); Fan and Fan (2008); Cai and Zhang (2019); Chen and Sun (2022) for $L>2$. However, it isn’t always reasonable to assume that such directions are sparse and the procedure is less attractive compared to regression-based approaches discussed below shortly. Witten and Tibshirani (2011) alternatively use the fact that the Bayes rule (1.2) is equivalent with Fisher’s discriminant rule
@@
$\begin{split}F_{k}:=\operatorname*{arg\,max}_{\beta\in\mathbb{R}^{p}}~{}\beta^{\top}\Sigma_{b}\beta\quad\text{subject to}&~{}\beta^{\top}\Sigma_{w}\beta=1,~{}\beta^{\top}\Sigma_{w}F_{i}=0,~{}~{}\forall~{}i<k.\end{split}$ (1.6)
Witten and Tibshirani (2011) develop methodology tailored to the high-dimensional scenarios with a diagonal matrix $\Sigma_{w}$ and sparse columns of $F$. However, the implicit formulation (1.6) of the matrix $F$ makes it extremely difficult to obtain the global solution and to derive theoretical properties of the regularized estimator and the resulting classifier.
More closely related to our approach is optimal scoring which dates back to De Leeuw et al. (1976); Young et al. (1978) and regains attention in the LDA setting, see Hastie et al. (1995, 1994); Clemmensen et al. (2011); Gaynanova (2020) and the references therein. By writing the response matrix and feature matrix as
${\mathbf{Y}}=(Y_{1},\ldots,Y_{n})^{\top}\in\{0,1\}^{n\times L},\quad{\mathbf{X}}=(X_{1},\ldots,X_{n})^{\top}\in\mathbb{R}^{n\times p},$ (1.7)
@@
$\begin{split}\min_{\ \theta\in\mathbb{R}^{L},\ \beta\in\mathbb{R}^{p}}&~{}\|{\mathbf{Y}}\theta-{\mathbf{X}}\beta\|_{2}^{2}\quad\text{s.t.}~{}\theta^{\top}{\mathbf{Y}}^{\top}{\mathbf{Y}}\theta=n,~{}\theta^{\top}{\mathbf{Y}}^{\top}{\mathbf{Y}}\widehat{\theta}_{i}=0,~{}\forall~{}i<k.\end{split}$ (1.8)
@@
Here $\widehat{\theta}_{i}$ is the solution to (1.8) in the $i$th iteration. In the low-dimensional case ($p<n$), Hastie et al. (1995) show that the solutions $\widehat{\beta}_{1},\ldots,\widehat{\beta}_{K}$ from (1.8) can be computed from canonical correlation analysis (CCA), and are further related, via CCA, with the matrix $F$ obtained from the sample analogue of (1.6). As a result, classification can be based on the dimension reduction directions $\widehat{\beta}_{1},\ldots,\widehat{\beta}_{K}$. In the high-dimensional setting, regularization in (1.8) is needed. For instance, Hastie et al. (1995) consider
@@
$\begin{split}\min_{\theta\in\mathbb{R}^{L},\ \beta\in\mathbb{R}^{p}}&~{}\|{\mathbf{Y}}\theta-{\mathbf{X}}\beta\|_{2}^{2}+\beta^{\top}\Omega\beta\quad\text{s.t.}~{}\theta^{\top}{\mathbf{Y}}^{\top}{\mathbf{Y}}\theta=n,~{}\theta^{\top}{\mathbf{Y}}^{\top}{\mathbf{Y}}\widehat{\theta}_{i}=0,~{}\forall~{}i<k,\end{split}$ (1.9)
@@
for a chosen $p\times p$ symmetric, positive definite matrix $\Omega$, and show that the global solution to (1.9) can be computed. A possible choice of $\Omega=\lambda\mathbf{I}_{p}$ for some $\lambda>0$ leads to the ridge penalty (Campbell, 1980; Friedman, 1989). To accommodate more general regularization, Gaynanova (2020) studies the matrix formulation of (1.9),
$\begin{split}\min_{\Theta,\ B}&~{}\|{\mathbf{Y}}\Theta-{\mathbf{X}}B\|_{F}^{2}+\text{pen}(B)\quad\text{s.t.}~{}\Theta^{\top}{\mathbf{Y}}^{\top}{\mathbf{Y}}\Theta=n\mathbf{I}_{K},\quad\Theta^{\top}{\mathbf{Y}}^{\top}{\mathbf{Y}}1_{n}=0_{K},\end{split}$ (1.10)
@@
and proves that the global solution to (1.10) can be computed, provided pen$(B)$ satisfies $\text{pen}(BR)=\text{pen}(B)$ for any orthogonal $K\times K$ matrix $R$. Gaynanova (2020) focuses on the group-lasso penalty $\|B\|_{1,2}=\sum_{j=1}^{p}\|B_{j\cdot}\|_{2}$ and analyzes the resulting classifier. However, it is not clear how the analysis can be extended to other penalized estimation. Moreover, the requirement that pen($B$) is orthogonal invariant excludes the familiar lasso penalty $\|B\|_{1}=\sum_{j,k}|B_{jk}|$ as well as its extensions such as the elastic net penalty. That being said, for penalties that are not orthogonal invariant, there is no guarantee of computing the global solution to (1.10), let alone any theoretical property of the resulting estimator.
@@
Key to our approach is the following reformulation of the problem. We show in Lemma 1 of Section 2 that the discriminant directional matrix under model (1.1)
$B^{*}=\Sigma_{w}^{-1}{M}\in\mathbb{R}^{p\times L}$ (1.11)
@@
$B:=\Sigma^{-1}\Sigma_{XY}:=[{\rm Cov}(X)]^{-1}{\rm Cov}(X,Y)\in\mathbb{R}^{p\times L}$ (1.12)
@@
$B^{*}$ $=BH^{-1}$ (1.13)
for some invertible $L\times L$ matrix $H$ with explicit expression. This connection is novel and distinguishes our approach from the current literature. Although the fact that columns of $B$ and $B^{*}$ share the same span is known (see, for instance, Hastie et al. (1995); Ye (2007); Lee and Kim (2015); Nie et al. (2022) and the references therein), their closed-form connection given by (1.13) has not been discovered, and a thorough theoretical investigation on the corresponding classifier has not been conducted. To the best of our knowledge, only in the binary case ($L=2$), Mai et al. (2012) leverages sparse linear regression between an appropriately encoded label vector and the feature matrix ${\mathbf{X}}$ for classification and proves consistency. Extension of such procedure to handle multi-class responses with high-dimensional features has been a longstanding open problem. In view of (1.13), we propose to estimate the discriminant direction matrix $B^{*}$ by first estimating the regression matrix $B$ as
$\widehat{B}=\operatorname*{arg\,min}_{B\in\mathbb{R}^{p\times L}}~{}{1\over n}\|{\mathbf{Y}}-{\mathbf{X}}B\|_{F}^{2}+\text{pen}(B),$ (1.14)
and then estimating $H^{-1}$ based on $\widehat{B}$ via its explicit expression in Lemma 1. The advantages of our approach as well as our main contributions are three-fold:
@@
    Akin to optimal scoring, our approach in (1.14) is regression-based, which, as pointed out in Hastie et al. (1995, 1994), is often easier to compute and more amenable to extensions to regularized and non-parametric regression as well as model selection than canonical correlation analysis, Fisher’s discriminant rule and the aforementioned procedures based on approximately solving the equation $\Sigma_{w}B^{*}=M$ for $B^{*}$. We can also easily tap into the large literature on estimation of the high-dimensional regression matrix $B\in\mathbb{R}^{p\times L}$ in (1.14).
@@
    After computing $\widehat{B}$ from (1.14), we only need to estimate the inverse of an $L\times L$ matrix to estimate the discriminant direction matrix $B^{*}$ rather than that of a $p\times p$ matrix $\Sigma_{w}$ as in the classical LDA context. Our approach in Section 2.2 uses the explicit form of $H$ in Lemma 1 and enjoys computational advantage as well as numerical stability, compared to the existing approaches. In Section 2.3, we further offer an alternative way to estimate the Bayes rule that performs Fisher’s discriminant analysis on low-dimensional transformations $\widehat{B}^{\top}x\in\mathbb{R}^{L}$ of the original features $x\in\mathbb{R}^{p}$. This leads to a practically useful method, which could extract fewer discriminant directions than those in $B^{*}$, for both downstream analysis and visualization.
@@
    We are able to quantify the errors of estimating both $B$ and $B^{*}$ as well as to provide theoretical guarantees for the resulting classifier, thanks to (1.13). Concretely, in Section 3.1 and Section 3.2 we provide a general strategy for analyzing the excess misclassification risk of the proposed classifier, which is valid for a generic estimator $\widehat{B}$ of $B$. While various choices of $\widehat{B}$ are feasible in our approach, more specifically in (1.14), we apply the general result to two popular estimators: the lasso estimator in Section 3.3 and the reduced rank estimator in Section 3.4. The former also contrasts our approach with Cai and Zhang (2019) and Witten and Tibshirani (2011) which assume sparsity of $B^{*}$ and $F$, respectively. The latter is particularly suitable when the rank of the discriminant direction matrix $B^{*}$ is low and the number of classes $L$ is moderate / large. For instance, our earlier work Bing and Wegkamp (2023) considers a multi-class classification setting with high-dimensional features $X\in\mathbb{R}^{p}$ that follow a factor model $X=AZ+W$ with unobserved (latent) features $Z\in\mathbb{R}^{r}$ with $r<p$ and random noise $W$, that is independent of both $Z$ and the label $Y$. In this case, the discriminant direction matrix $B^{*}$ has rank no greater than $r$ when $r<L$. The procedure in Bing and Wegkamp (2023), however, is different than the one proposed here. We refer to Remark 1 in Section 2.2 for more discussion.
Finally, we provide a thorough simulation study in Section 4 to corroborate our theoretical findings. In Section 5, we report our findings in several real data studies. The Appendix contains most of the proofs and additional simulation study.
Notation.
For any numbers $a,b\in\mathbb{R}$, we write $a\vee b=\max\{a,b\}$. For any vector $v$ and $1\leq q\leq\infty$, we use $\|v\|_{q}$ to denote the standard $\ell_{q}$-norm. For any positive integer $d$, we write $[d]:=\{1,\ldots,d\}$. For any matrix $Q\in\mathbb{R}^{d_{1}\times d_{2}}$, any $i\in[d_{1}]$ and $j\in[d_{2}]$, we write $Q_{i\cdot}$ for its $i$th row and $Q_{j}$ for its $j$th column. For norms, we write $\|Q\|_{\infty}=\max_{i,j}|Q_{ij}|$, $\|Q\|_{F}^{2}=\sum_{i,j}Q_{ij}^{2}$, $\|Q\|_{1}=\sum_{i,j}|Q_{ij}|$, $\|Q\|_{1,2}=\sum_{i}\|Q_{i\cdot}\|_{2}$ and $\|Q\|_{\rm op}=\sup_{v:\|v\|_{2}=1}\|Qv\|_{2}$. For any symmetric, semi-positive definite matrix $Q\in\mathbb{R}^{d\times d}$, we use $\lambda_{1}(Q),\lambda_{2}(Q),\ldots,\lambda_{d}(Q)$ to denote its eigenvalues in non-increasing order. For any sequences $a_{n}$ and $b_{n}$, we write $a_{n}\lesssim b_{n}$ if there exists some constant $C>0$ such that $a_{n}\leq Cb_{n}$, and $a_{n}\asymp b_{n}$ if $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$.